# Enumeration of Elliptic Curves via Elliptic Gromov-Witten Invariants of Four Dimensional Projective Fano Hypersurfaces

### v1

@@ Abstract
In this note, we propose a conjecture that clarifies the relationship between the number of degree $d$ elliptic curves in complex four-dimensional projective Fano hypersurfaces and their degree $d$ elliptic Gromov-Witten (GW) invariants. The elliptic GW invariants are computed using the elliptic virtual structure constants proposed in our previous works.
@@ 1 Introduction
Now, let us turn to the case of genus $1$ GW invariants. The genus $1$ GW invariants of projective Fano surfaces (complex two-dimensional manifolds) are non-negative integers, at least for lower degrees. However, in the Calabi-Yau case, we know that the elliptic GW invariants include not only contributions from multiple cover maps from a genus $1$ curve but also degenerate contributions associated with genus 0 GW invariants. Roughly speaking, the degenerate contributions arise from the boundary loci of the moduli space of stable maps from genus $1$ stable curves, where the stable curves degenerate into nodal curves containing a genus 0 component. At these loci, the stable curves can be mapped into rational curves in the target manifold $X$, and the contributions are related to the genus 0 GW invariants of $X$. It is well-known that degenerate contributions also appear in the case of complex three-dimensional Fano manifolds. Heuristically, this phenomenon is expected from the dimensional counting of the virtual dimension of the moduli space of stable maps. In the genus $1$ case, the virtual dimension of the moduli space of stable maps from stable curves to complex three-dimensional projective manifolds with fixed degree $d$ is independent of the genus of the stable curves. Hence, we can speculate that the genus $1$ GW invariants may contain degenerate contributions because the operator insertions required to balance the positive virtual dimension of the moduli space do not depend on the genus. A natural question thus arises: **”Do degenerate contributions appear in the case of a projective Fano manifold whose complex dimension is greater than 3?”**
@@
In this note, motivated by the above question, we investigate the relationship between the elliptic GW invariants of four-dimensional projective Fano manifolds and the number of elliptic curves. In the case of complex four-dimensional Kähler manifolds, the virtual dimension of the moduli space of stable maps from genus 0 stable curves is one complex dimension greater than that from genus $1$ stable curves. Therefore, it is not clear whether degeneration contributions appear, based on the naive speculation used in the preceding discussion. In the Calabi-Yau case, the work [13] by Klemm and Pandharipande shows that degenerate contributions do indeed appear, and they are related to genus 0 GW invariants in a more complicated way than in the case of complex three-dimensional Calabi-Yau manifolds. We must note that their results rely on the computation of genus $1$ GW invariants using the BCOV conjecture [2]. On the other hand, in [7], we proposed a conjecture that enables a kind of B-model computation (from the perspective of mirror symmetry) of genus $1$ GW invariants for projective hypersurfaces of arbitrary dimension. We therefore used our conjecture to evaluate genus $1$ GW invariants of four-dimensional projective Fano hypersurfaces. In the Fano case, the Virasoro conjecture [4] and Getzler’s equation [6] are also effective for this purpose. When applying these methods to evaluate genus $1$ invariants, one must use information from genus 0 GW invariants with insertions of operators derived from primitive cohomology classes of the hypersurfaces. As was done in [5], these genus 0 invariants were computed in the case of a cubic hypersurface in $CP^{4}$ using the associativity equation, since it has primitive classes only in the bidegree $(2,1)$ and $(1,2)$ sectors. However, in the four-dimensional case, a projective Fano hypersurface can have primitive classes in the bidegree $(3,1)$, $(2,2)$, and $(1,3)$ sectors. Therefore, we must overcome technical obstacles to compute genus 0 invariants with these operator insertions using known techniques. Moreover, as we experienced in the computation of genus $1$ invariants of the cubic hypersurface using the Virasoro conjecture [5], the method requires outrageously complicated and lengthy computations. In contrast, our conjecture does not require any information regarding genus 0 GW invariants with insertions of primitive classes and can be easily automated using computer software such as Maple and Mathematica.
@@
Acknowledgment
We thank Prof. H. Suzuki for valuable discussions. Our research is partially supported by JSPS Grant No. 22K03289. The research of M. J. is also partially supported by JSPS Grant No. 24H00182.
@@ 2 Definitions and Our Main Conjecture
$\langle{\cal O}_{h}\prod_{a=0}^{N-2}({\cal O}_{h^{a}})^{m_{a}}\rangle_{g,d}=d\langle{\cal O}_{h}\prod_{a=0}^{N-2}({\cal O}_{h^{a}})^{m_{a}}\rangle_{g,d}\quad(d\geq 1),$ (2.4)
@@ Appendix A Enumeration of Degree $3$ Elliptic Curves in $M_{6}^{k}$ ($k=1,2,3,4,5$)
We find that $\tilde{E}_{3,a,b,c}$ indeed coincides with $E_{3,a,b,c}$.

### v4

@@ Abstract
In this paper, we propose a conjecture that clarifies the relationship between the number of degree $d$ elliptic curves in complex four-dimensional projective Fano hypersurfaces and their degree $d$ elliptic Gromov-Witten (GW) invariants. The elliptic GW invariants are computed using the elliptic virtual structure constants proposed in our previous works [^1] Keywords; enumerative geometry; gromov-witten invariants; mirror symmetry; elliptic curve; projective Fano hypersurface. 2020 MATHEMATICS SUBJECT CLASSIFICATION; 14N10, 14N35. .
@@ 1 Introduction
Now, let us turn to the case of genus $1$ GW invariants. The genus $1$ GW invariants of projective Fano surfaces (complex two-dimensional manifolds) are non-negative integers, at least for lower degrees. However, in the Calabi-Yau case, we know that the elliptic GW invariants include not only contributions from multiple cover maps from a genus $1$ curve but also degenerate contributions associated with genus 0 GW invariants [2, 15]. Roughly speaking, the degenerate contributions arise from the boundary loci of the moduli space of stable maps from genus $1$ stable curves, where the stable curves degenerate into nodal curves containing a genus 0 component. At these loci, the stable curves can be mapped into rational curves in the target manifold $X$, and the contributions are related to the genus 0 GW invariants of $X$. It is well-known that degenerate contributions also appear in the case of complex three-dimensional Fano manifolds. Heuristically, this phenomenon is expected from the dimensional counting of the virtual dimension of the moduli space of stable maps. The virtual dimension of the moduli space of stable maps from stable curves to complex three-dimensional projective manifolds with fixed degree $d$ is independent of the genus of the stable curves. Hence, we can speculate that the genus $1$ GW invariants may contain degenerate contributions because the operator insertions required to balance the positive virtual dimension of the moduli space do not depend on the genus. A natural question thus arises: **”Do degenerate contributions appear in the case of a projective Fano manifold whose complex dimension is greater than 3?”**
@@
In this paper, motivated by the above question, we investigate the relationship between the elliptic GW invariants of four-dimensional projective Fano manifolds and the number of elliptic curves. In the case of complex four-dimensional Kähler manifolds, the virtual dimension of the moduli space of stable maps from genus 0 stable curves is one complex dimension greater than that from genus $1$ stable curves. Therefore, it is not clear whether degeneration contributions appear, based on the naive speculation used in the preceding discussion. In the Calabi-Yau case, the work [14] by Klemm and Pandharipande shows that degenerate contributions do indeed appear, and they are related to genus 0 GW invariants in a more complicated way than in the case of complex three-dimensional Calabi-Yau manifolds. We must note that their results rely on the computation of genus $1$ GW invariants using the BCOV conjecture [2]. On the other hand, in [7], we proposed a conjecture that enables a kind of B-model computation (from the perspective of mirror symmetry) of genus $1$ GW invariants for projective hypersurfaces of arbitrary dimension. We therefore used our conjecture to evaluate genus $1$ GW invariants of four-dimensional projective Fano hypersurfaces. In the Fano case, the Virasoro conjecture [4] and Getzler’s equation [6] are also effective for this purpose. When applying these methods to evaluate genus $1$ invariants, one must use information from genus 0 GW invariants with insertions of operators derived from primitive cohomology classes of the hypersurfaces. As was done in [5], these genus 0 invariants were computed in the case of a cubic hypersurface in $CP^{4}$ using the associativity equation, since it has primitive classes only in the bidegree $(2,1)$ and $(1,2)$ sectors. However, in the four-dimensional case, a projective Fano hypersurface can have primitive classes in the bidegree $(3,1)$, $(2,2)$, and $(1,3)$ sectors. Therefore, we must overcome technical obstacles to compute genus 0 invariants with these operator insertions using known techniques. Moreover, as we experienced in the computation of genus $1$ invariants of the cubic hypersurface using the Virasoro conjecture [5], the method requires outrageously complicated and lengthy computations. In contrast, our conjecture does not require any information regarding genus 0 GW invariants with insertions of primitive classes and can be easily automated using computer software such as Maple and Mathematica.
@@ 2 Definitions and Our Main Conjecture
$\langle{\cal O}_{h}\prod_{a=0}^{N-2}({\cal O}_{h^{a}})^{m_{a}}\rangle_{g,d}=d\langle\prod_{a=0}^{N-2}({\cal O}_{h^{a}})^{m_{a}}\rangle_{g,d}\quad(d\geq 1),$ (2.4)
@@ Appendix A Enumeration of Degree $3$ Elliptic Curves in $M_{6}^{k}$ ($k=1,2,3,4,5$)
We find that $\tilde{E}_{3,a,b,c}$ indeed coincides with $E_{3,a,b,c}$  [^2] A PDF copy of the Maple worksheet used for computing $\tilde{E}_{3,a,b,c}$ of $M_{6}^{5}$ is available on M. Jinzenji’s ResearchGate homepage [9].
Acknowledgment
We would like to thank Prof. H. Suzuki for valuable discussions.
Funding
Our research is partially supported by JSPS Grant No. 22K03289. The research of M. J. is also partially supported by JSPS Grant No. 24H00182.
Disclosure Statements on Competing Interests
We have no relevant financial or non-financial ompeting interests to declare.